how many gallons of 5% alcohol solution must be mixed with 90 gallons of 1% solution to obtain a 2% solution

Answer by dkppathak(439)

let x gallon 5% solution added in 90 gallon of 1% solutions
alcohol available in 5 % solution =5x/100
alcohol available in 1% 90 gallon =90/100 gallon
after mixing alcohol available in 2% solution=(90+x)2/100
as per given conditions
5x/100+90/100=2(90+x)/100
5x+90=180+2x
5x-2x=180-90
3x=90
x=90/3
x=30 gallon
answer 30 gallon of 5% alcohol is to be added in 90 gallon 1% solution to get 2%alcohol solution